# Regularity and temperature of stationary black hole event horizons

### v1

@@ 2 Existence of the principal bundles
Proof.
By [35, Prop. 4.82] the set at which strong causality holds is open, thus there is an open neighborhood $U$ of $H^{+}$ at which strong causality holds. By Prop. 2.11 $k\neq 0$ on $H^{+}$ hence, by continuity, there is an open neighborhood $U^{\prime}$ of $H^{+}$ over which $k\neq 0$. The open neighborhood of $H^{+}$, $R:=\phi(U\cap U^{\prime})$, is such that $k\neq 0$ on it and strong causality holds on it. By Prop. 2.14 we can find an invariant open neighborhood $Z$ as in that result. Let $W=R\cap Z$ and $V:=W\cap I^{+}(M_{end})$.
Suppose the action $\phi$ is not free on $V$. Then we can find $p\in V$ and $t\neq 0$ such that $\phi_{t}(p)=p$, which implies $p=\phi_{-t}(p)$ and so for every $m\in\mathbb{Z}$, $\phi_{mt}(p)=p$. But for $q\in M_{end}$, $p\in I^{+}(r)$ for some $r\in\gamma_{q}$, which implies $p\in I^{+}(\phi_{mt}(r))$ for every $m\in\mathbb{Z}$ and hence $\gamma_{q}\subset I^{-}(p)$, in contradiction with the property of Prop. 2.14 for $K=\{p\}$. This shows that the action $\phi$ is free.
Let $r\in M_{end}$ and let $K_{1},K_{2}\subset V$ be compact subsets. By Cor. 2.7 for each $p\in K_{1}$ we can find some $q\in\gamma_{r}$ such that $p\in I^{+}(q)$, thus, passing to a finite subcovering, we see that we can choose $x\in\gamma_{r}$ such that $K_{1}\subset I^{+}(x)$. Moreover, we know from Prop. 2.14, and from $K_{2}\subset Z$ that there is $y\in\gamma_{r}$ such that $K_{2}\cap\overline{I^{+}(y)}=\emptyset$, so that the same is true for every $y^{\prime}\geq y$, $y^{\prime}\in\gamma_{r}$. We can choose $y$ so that $y\gg x$. Let $\tau>0$ be such that $\phi_{\tau}(x)=y$. Then it cannot be $\phi_{t}(K_{1})\cap K_{2}\neq\emptyset$ for any $t\geq\tau$.
@@
As $t$ such that $\phi_{t}(K_{1})\cap K_{2}\neq\emptyset$ is upper bounded, so is $s$ such that $K_{1}\cap\phi_{s}(K_{2})\neq\emptyset$. But since the latter equation is equivalent to $\phi_{-s}(K_{1})\cap K_{2}\neq\emptyset$, we conclude that the $t$ such that $\phi_{t}(K_{1})\cap K_{2}\neq\emptyset$ are bounded. The map $\phi:(t,x)\mapsto(\phi_{t}(x),x)$ is continuous, thus $\phi^{-1}(K_{2}\times K_{1})$ is closed, but $\phi^{-1}(K_{2}\times K_{1})\subset I\times K_{1}$ with $I$ compact interval, which proves properness. ∎
@@
From the standard result [2, Thm. 3.34] [15, Thm. 1.11.4 55] [31, Thm. 21.10] we obtain (by manifold we understand Hausdorff manifold)
Corollary 2.16.
Suppose that $(M,g)$ is strongly causal at $H^{+}$. The quotient $V/\mathbb{R}$ is a manifold and so $\pi:V\to V/\mathbb{R}$ is a trivial principal bundle with structure group $(\mathbb{R},+)$. As a consequence, $H^{+}$ is diffeomorphic to the product ×$\mathbb{R}\times S$, with $S:=\pi(H^{+})\subset B:=V/\mathbb{R}$, where $S$ is closed subset of $V/\mathbb{R}$, and where the orbits are the whole $\mathbb{R}$-fibers.
@@
Since the projection is a quotient map and $H^{+}=\pi^{-1}(S)$ is closed, it follows that $S$ is closed. ∎
Corollary 2.17.
Suppose that $(M,g)$ is strongly causal at $H^{+}$. There is a smooth codimension one hypersurface $\Sigma$ in $V$ which intersects exactly once every Killing orbit (hence those belonging to $H^{+}$) and which is diffeomorphic to $B$. Its intersection $\sigma=\Sigma\cap H^{+}$ is diffeomorphic to $S$ and hence has the same number of components of $H^{+}$. By removing it $H^{+}\backslash\sigma$ gets twice the original number of components.
@@
Proof.
All the results are consequence of the $\mathbb{R}$-bundle being trivial. The hypersurface $\Sigma$ is just a level set $t=cost$ from the trivializing diffeomorphism. ∎
Proposition 2.18.
@@
Let $r\in M_{end}$ so that, by the construction of $V$, the property of Prop. 2.14 holds, namely for every $K\subset V$ we can find $r^{\prime}\in\gamma_{r}$ such that $K\cap\overline{I^{+}(r^{\prime})}=\emptyset$.
Suppose that there is $n\neq 0$ and $p\in V$ such that $\phi_{n\tau}(p)=p$. Since $I^{+}(M_{end})=I^{+}(\gamma_{r})$ we know that there is some $z\in\gamma_{r}$ such that $p\in I^{+}(z)$ and hence for every $k\in\mathbb{R}$, $p\in I^{+}(\phi_{kn\tau}(z))$. This implies $I^{-}(p)\supset\phi(z)=\gamma_{r}$. But by Prop. 2.14 this is not possible, just set $K=\{p\}$. ∎
Corollary 2.19.
Let $(M,g)$ be strongly causal at $H^{+}$. Let $\tau>0$ and consider the action $\psi$ as in Prop. 2.18. The quotient $V/\mathbb{Z}$ is a manifold, and there is a normal covering map $c:V\to V/\mathbb{Z}$. Furthermore, there is a trivial $S^{1}$-principal bundle $\hat{\pi}:V/\mathbb{Z}\to V/\mathbb{R}$ with $\pi=\hat{\pi}\circ c$.
Proof.
The group $\mathbb{Z}$ endowed with the discrete topology is, being countably infinite, a discrete Lie group. We have already established that its action is proper and free. By [31, Thm. 21.13] there exists a normal covering map $c:V\to V/\mathbb{Z}$.
@@
The last bundle is trivial because, if $s:B\to V$, $B=V/\mathbb{R}$, is a smooth section for $\pi:V\to B$, then $c\circ s:B\to V/\mathbb{Z}$ is a section for $\hat{\pi}:V/\mathbb{Z}\to B$. ∎
@@ 3 Proving smoothness
The next paragraphs and Thm. 3.1 are given to provide characterization (b) in Def. 3.2 of ‘horizon with compact section’ but can be skipped on first reading.
@@
The next result clarifies what it means for the horizon to have compact space sections. It does not demand a compact set to intersect all the generators or all the Killing orbits once, as both requests are strong and should rather be deduced from weaker assumptions. Our disconnection assumption does not mention generators (compare [7, Sec. 4.1]).
Theorem 3.1.
@@
The next definition should not be confused with the definition given in [7, Sec. 4.1]. Only in the next section we shall prove the existence of a compact subset of $H^{+}$ intersected by every generator precisely once, a result which is assumed in [7] and which is basically a $C^{1}$ assumption on $H^{+}$.
Definition 3.2.
@@
Proof of the equivalence.
(a)$\Rightarrow$ (b). In the trivialization just let $K$ be any level set of $t$. (b)$\Rightarrow$ (a). This is Theorem 3.1. ∎
Theorem 3.3.
Let $(M,g)$ be a spacetime endowed with a $C^{3}$ metric that satisfies the null convergence condition. Let $k$ be a Killing field whose orbits are complete. Let $\Sigma_{end}$ be an acausal hypersurface, possibly with edge, over which $k$ is timelike. Suppose that the horizon $H^{+}:=\partial I^{-}(M_{end})\cap I^{+}(M_{end})$, is connected, has compact projection (cf. Def. 3.2) and that strong causality holds on it. If $\theta\geq 0$ on $H^{+}$ in the sense of support functions (implied by variants of asymptotic flatness condition, see discussion below) then $H^{+}$ is a totally geodesic future null hypersurface as smooth as the metric (e.g. smooth/analytic if the metric is smooth/analytic).
@@
In the context of spacetimes admitting a conformal completion with asymptotic infinity $\mathscr{I}^{+}$, the condition $\star$ can be proved under some reasonable causality conditions. For instance, Hawking deduced it from stronger but physically motivated conditions on the asymptotic structure, and in particular from the assumption of asymptotic predictability (weak cosmic censorship): $\mathscr{I}^{+}\subset\overline{D^{+}(S)}$ where $S$ is a partial Cauchy hypersurface and the closure is in the topology of $\bar{M}$. The reader is referred to [23, 8] for a discussion of the reasonability of $\star$. In the regularity result by Chruściel and Costa [7, Thm. 4.11] it is also present though framed again using $\mathscr{I}^{+}$, see also [7, Thm. 4.10]. The formulation $\star$ allows typically for shorter and less technical presentations.
As Hawking showed, $\star$ allows one to prove that the expansion is positive on the horizon [23, Lemma 9.2.2], a fact which ultimately leads to the proof of the second law of black hole thermodynamics which states that the area of black holes is non-decreasing (see [8, 34] for a proof without smoothness assumptions). The fact that Hawking’s argument on the positivity of $\theta$ can be adapted to the non-smooth case is non-trivial and was proved in [8, Theor. 4.1]. These authors still work in the geometry of spacetime admitting a suitable conformal completion. This requires more assumptions, though the framework is compatible with that adopted in this work, in which the horizon is defined via the boundary of $I^{-}(M_{end})$. We adapt the result [8, Theor. 4.1] to the present framework as follows.
Theorem 3.5.
@@
A proof can be obtained from the sketch of proof in [34, Thm. 22] (see also the original reference [8, Thm. 4.1]), with some trivial replacements such as $I^{-}(\mathscr{I}^{+},\bar{M})\to I^{-}(M_{end})$. Of course, the inequality $\theta\geq 0$ is understood in a support function sense, as $H^{+}$ is a priori non-differentiable.
@@
Proof of Theorem 3.3.
Under the quotient of Corollary 2.19 the horizon $H^{+}$ projects to a compact $C^{0}$ future null hypersurface $\hat{H}$ which is topologically a product. Moreover, we have $\theta\geq 0$ on $\hat{H}$, thus by the result in [30, Thm. 1.43] [34, Thm. 18] it follows that $\hat{H}$ and hence $H^{+}$ is a totally geodesic lightlike hypersurface as regular as the metric (the Cauchy horizon condition in these theorems is used to infer future completeness of the generators from which $\theta\geq 0$ is inferred, thus for our purposes the Cauchy horizon condition can be replaced by $\theta\geq 0$, indeed [34, Thm. 13] implies that $\theta=0$, $\mu^{s}_{ij}=0$ and [34, Thm. 17] implies that it is as smooth as the metric). Notice that although through each point of $H^{+}$ passes a unique generator, the generators of $H^{+}$ need not be past-inextendible (in some examples they are past incomplete and can be extended escaping $H^{+}$). ∎
Remark 3.6.
We get smoothness starting from $\star$, and without passing through conclusions about the completeness of the generators. One could try to prove the dichotomy non-degenerate/degenerate (i.e. all generators are future complete (or all past complete), or all incomplete) of compact horizons [37, 38, 45, 20] stay in the future complete case, infer $\theta\geq 0$ from the achronality of the horizon and then get the smoothness of the horizon with a proof analogous to the above, and without using $\star$. The problem is that available proofs of the mentioned dichotomy apply to compact horizons that are already known to be smooth, so one would have first to extend the ribbon argument and other analytical techniques to the non-smooth horizon setting, a strategy which is not entirely clear could be successfully pursued given the technical difficulties involved.
@@ 4 Existence of cross-sections
Proof.
Let $\gamma:I\to H^{+}$ be a generator, and assume that there is $s\in I$ such that for some $r>0$, $\phi_{r}(\gamma(s))\leq\gamma(s)$. It follows that for every $\0n\in\mathbb{N}\backslash\{0\}$, $\phi_{nr}(\gamma(s))\leq\gamma(s)$. Let $q\in M_{end}$ be such that $\gamma(s)\notin I^{+}(q)$. Let $\gamma_{q}=\phi(q)$, $q=\gamma_{q}(0)$. We know that there is some $p=\gamma_{q}(a)$, $a<0$, such that $\gamma(s)\in I^{+}(p)$, but for $n$ so large that $a+nr>0$ we have $\phi_{nr}(\gamma(s))\in I^{+}(\phi_{nr}(p))\subset I^{+}(q)$, and hence $\gamma(s)\in I^{+}(q)$, a contradiction. ∎
In this section we shall tacitly make the following assumption:
strongly causality holds on $H^{+}$ and $S=\pi(H^{+})$ is compact.
@@
We recall that a cross-section in the sense of [7] is a topological submanifold of $H^{+}$ that intersects all generators exactly once. We shall indeed prove that cross-sections exist (in [7] this was an assumption) by taking advantage of the existence of the trivial principal bundle established in the previous section. Note that contrary to [7] we do not use the existence of cross-sections to obtain smoothness results, the logical order in this work is reversed ($C^{1}$ regularity is used to obtain cross-sections, see Lemma 4.4 below).
@@
The next result uses ideas in [16, Remark 2.2], see also [12]. In that remark they used without mention a causal simplicity assumption at the end of their argument (also beware that elsewhere they assume smoothness of the horizon, which, in any case, we proved without using this type of results). We are able to considerably weaken the causality condition.
Lemma 4.2.
@@
Let $r\in M_{end}$ then every Killing orbit of $H^{+}$ intersects $Y:=\partial I^{+}(r)\cap H^{+}$. Moreover, in the forward direction it is bound to enter $I^{+}(r)$ and to remain in it once entered. Similarly, in the backward direction, it is bound to escape $\overline{I^{+}(r)}$ and to remain outside it once escaped. Thus the orbit intersects $Y$ in a compact segment (possibly a point).
@@
If past reflectivity holds in an open set containing $\overline{I^{+}(M_{end})}\cap I^{+}(M_{end})$ then the intersection is just one single point. In this case $Y$ and $S$ are homeomorphic.
@@
The curious and interesting role of past reflectivity in black hole physics has been pointed out in [36]. We recall that it is implied by causal continuity and hence by global hyperbolicity. It is also implied by the existence of a continuous Lorentzian distance function [3] [35, Prop. 5.2]. Finally, it is implied by the existence of a timelike Killing field which is complete in the past direction [13] [35, Thm. 4.10]. As a consequence, past reflectivity holds on $M_{end}$, however, a priori it might not extend up to the horizon.
@@
It is worth recalling that past reflectivity and past distinction imply stable causality [35, Thm. 4.111] and that we are imposing strong causality in a neighborhood of $H^{+}$. So imposing the property of past reflectivity implies the existence of a time function in a neighborhood of the horizon.
Proof.
@@
Let $p\in H^{+}$ then as $H^{+}\subset I^{+}(\phi(r))$, we can find $\tau>0$ such that $p\in I^{+}(\phi_{-\tau}(r))$ (remember that $k$ is timelike in $M_{end}$ but not necessarily near the horizon), and hence $\phi_{\tau}(p)\in I^{+}(r)$. Moreover, by Prop. 2.13 we can find $s>0$ such that $p\notin\overline{I^{+}(\phi_{s}(r))}$ thus $\phi_{-s}(p)\notin\overline{I^{+}(r)}$. Thus the Killing orbit $x(t)=\phi_{t}(p)\subset H^{+}$ intersects $\partial I^{+}(r)$ and in the forward direction it enters $I^{+}(r)$ while in the past direction it escapes $\overline{I^{+}(r)}$.
@@
Assume past reflectivity. If $q\in\overline{I^{+}(r)}\cap H^{+}$ then by past reflectivity $r\in\overline{I^{-}(q)}$ and for $\tau>0$, as $r\gg\phi_{-\tau}(r)$, $q\in I^{+}(\phi_{-\tau}(r))$, and by the isometry $\phi_{\tau}(q)\in I^{+}(r)$, namely any orbit can have at most one point in $\overline{I^{+}(r)}$ as any subsequent point belongs to $I^{+}(r)$. The bundle projection $\pi$ is continuous and its restriction to $Y$ sends bijectively $Y$ to $S$. But $Y$ is compact and $S$ is Hausdorff thus such restriction is a homeomorphism. ∎
Lemma 4.3.
Let $r\in M_{end}$, then the set $Y:=\partial I^{+}(r)\cap H^{+}$ is non-empty and compact.
Proof.

### v2

@@ 2 Existence of the principal bundles
Proof
By (minguzzi18b, , Prop. 4.82) the set at which strong causality holds is open, thus there is an open neighborhood $U$ of $H^{+}$ at which strong causality holds. By Prop. 6 $k\neq 0$ on $H^{+}$ hence, by continuity, there is an open neighborhood $U^{\prime}$ of $H^{+}$ over which $k\neq 0$. The open neighborhood of $H^{+}$, $R:=\phi(U\cap U^{\prime})$, is such that $k\neq 0$ on it and strong causality holds on it. By Prop. 8 we can find an invariant open neighborhood $Z$ as in that result. Let $W=R\cap Z$ and $V:=W\cap I^{+}(M_{end})$.
Suppose the action $\phi$ is not free on $V$. Then we can find $p\in V$ and $t\neq 0$ such that $\phi_{t}(p)=p$, which implies $p=\phi_{-t}(p)$ and so for every $m\in\mathbb{Z}$, $\phi_{mt}(p)=p$. But for $q\in M_{end}$, $p\in I^{+}(r)$ for some $r\in\gamma_{q}$, which implies $p\in I^{+}(\phi_{mt}(r))$ for every $m\in\mathbb{Z}$ and hence $\gamma_{q}\subset I^{-}(p)$, in contradiction with the property of Prop. 8 for $K=\{p\}$ (or Cor. 2). This shows that the action $\phi$ is free.
Let $r\in M_{end}$ and let $K_{1},K_{2}\subset V$ be compact subsets. By Cor. 1 for each $p\in K_{1}$ we can find some $q\in\gamma_{r}$ such that $p\in I^{+}(q)$, thus, passing to a finite subcovering, we see that we can choose $x\in\gamma_{r}$ such that $K_{1}\subset I^{+}(x)$. Moreover, we know from Prop. 8, and from $K_{2}\subset Z$ that there is $y\in\gamma_{r}$ such that $K_{2}\cap\overline{I^{+}(y)}=\emptyset$, so that the same is true for every $y^{\prime}\geq y$, $y^{\prime}\in\gamma_{r}$. We can choose $y$ so that $y\gg x$. Let $\tau>0$ be such that $\phi_{\tau}(x)=y$. Then it cannot be $\phi_{t}(K_{1})\cap K_{2}\neq\emptyset$ for any $t\geq\tau$.
@@
As $t$ such that $\phi_{t}(K_{1})\cap K_{2}\neq\emptyset$ is upper bounded, so is $s$ such that $K_{1}\cap\phi_{s}(K_{2})\neq\emptyset$. But since the latter equation is equivalent to $\phi_{-s}(K_{1})\cap K_{2}\neq\emptyset$, we conclude that the $t$ such that $\phi_{t}(K_{1})\cap K_{2}\neq\emptyset$ are bounded. The map $\phi:(t,x)\mapsto(\phi_{t}(x),x)$ is continuous, thus $\phi^{-1}(K_{2}\times K_{1})$ is closed, but $\phi^{-1}(K_{2}\times K_{1})\subset I\times K_{1}$ with $I$ compact interval, which proves properness. $\square$
@@
From the standard result (alexandrino15, , Thm. 3.34) (duistermaat00, , Thm. 1.11.4 55) (lee13, , Thm. 21.10) we obtain (by manifold we understand Hausdorff manifold)
Corollary 3
Suppose that $(M,g)$ is strongly causal at $H^{+}$. The quotient $V/\mathbb{R}$ is a manifold and so $\pi:V\to V/\mathbb{R}$ is a trivial principal bundle with structure group $(\mathbb{R},+)$. As a consequence, $H^{+}$ is diffeomorphic to the product $\mathbb{R}\times S$, with $S:=\pi(H^{+})\subset B:=V/\mathbb{R}$, where $S$ is closed subset of $V/\mathbb{R}$, and where the orbits are the whole $\mathbb{R}$-fibers.
@@
Since the projection is a quotient map and $H^{+}=\pi^{-1}(S)$ is closed, it follows that $S$ is closed. $\square$
Corollary 4
Suppose that $(M,g)$ is strongly causal at $H^{+}$. There is a smooth codimension one hypersurface $\Sigma$ in $V$ which intersects exactly once every Killing orbit (hence those belonging to $H^{+}$) and which is diffeomorphic to $B$. Its intersection $\sigma=\Sigma\cap H^{+}$ is diffeomorphic to $S$ and hence has the same number of components of $H^{+}$. By removing it, $H^{+}\backslash\sigma$ gets twice the original number of components.
@@
Proof
All the results are consequence of the $\mathbb{R}$-bundle being trivial. The hypersurface $\Sigma$ is just a level set $t=cost$ from the trivializing diffeomorphism. $\square$
Proposition 9
@@
Let $r\in M_{end}$ so that, by the construction of $V$, the property of Prop. 8 holds, namely for every $K\subset V$ we can find $r^{\prime}\in\gamma_{r}$ such that $K\cap\overline{I^{+}(r^{\prime})}=\emptyset$.
Suppose that there is $n\neq 0$ and $p\in V$ such that $\phi_{n\tau}(p)=p$. Since $I^{+}(M_{end})=I^{+}(\gamma_{r})$ we know that there is some $z\in\gamma_{r}$ such that $p\in I^{+}(z)$ and hence for every $k\in\mathbb{Z}$, $p\in I^{+}(\phi_{kn\tau}(z))$. This implies $I^{-}(p)\supset\phi(z)=\gamma_{r}$. But by Prop. 8 this is not possible, just set $K=\{p\}$. $\square$
Corollary 5
Let $(M,g)$ be strongly causal at $H^{+}$. Let $\tau>0$ and consider the action $\psi$ as in Prop. 9. The quotient $V/\mathbb{Z}$ is a manifold, and there is a normal covering map $c:V\to V/\mathbb{Z}$. Furthermore, there is a trivial $S^{1}$-principal bundle $\hat{\pi}:V/\mathbb{Z}\to V/\mathbb{R}$ with $\pi=\hat{\pi}\circ c$.
Proof
The group $\mathbb{Z}$ endowed with the discrete topology is, being countably infinite, a discrete Lie group. We have already established that its action is proper and free. By (lee13, , Thm. 21.13) there exists a normal covering map $c:V\to V/\mathbb{Z}$.
@@
The last bundle is trivial because, if $s:B\to V$, $B=V/\mathbb{R}$, is a smooth section for $\pi:V\to B$, then $c\circ s:B\to V/\mathbb{Z}$ is a section for $\hat{\pi}:V/\mathbb{Z}\to B$. $\square$
@@ 3 Proving smoothness
The next paragraphs and Thm. 3.1 are given to provide characterization (b) in Def. 2 of ‘horizon with compact section’ but can be skipped on first reading.
@@
The next result clarifies what it means for the horizon to have compact space sections. It does not demand a compact set to intersect all the generators or all the Killing orbits once, as both requests are strong and should rather be deduced from weaker assumptions. Our disconnection assumption does not mention generators (compare (chrusciel08, , Sec. 4.1)).
Theorem 3.1
@@
The next definition should not be confused with the definition given in (chrusciel08, , Sec. 4.1). Only in the next section we shall prove the existence of a compact subset of $H^{+}$ intersected by every generator precisely once, a result which is assumed in chrusciel08 and which is basically a $C^{1}$ assumption on $H^{+}$.
Definition 2
@@
Similar definitions apply to the connected components of $H^{+}$.
Proof (of the equivalence)
[…]
Theorem 3.2
Let $(M,g)$ be a spacetime endowed with a $C^{3}$ metric that satisfies the null convergence condition. Let $k$ be a complete Killing vector field. Let $\Sigma_{end}$ be an acausal hypersurface, possibly with edge, over which $k$ is timelike. Let $M_{end}:=\phi(\Sigma_{end})$ and suppose that the horizon $H^{+}:=\partial I^{-}(M_{end})\cap I^{+}(M_{end})$, is connected, has compact projection (cf. Def. 2) and that strong causality holds on it. If $\theta\geq 0$ on $H^{+}$ in the sense of support functions (implied by variants of asymptotic flatness condition, see discussion below) then $H^{+}$ is a totally geodesic future null hypersurface as regular as the metric (e.g. smooth/analytic if the metric is smooth/analytic). If $H^{+}$ is not connected the result holds on each connected component with compact projection provided strong causality and the condition $\theta\geq 0$ hold over such component.
@@
In the context of spacetimes admitting a conformal completion with asymptotic infinity $\mathscr{I}^{+}$, the condition $\star$ can be proved under some reasonable causality conditions. For instance, Hawking deduced it from stronger but physically motivated conditions on the asymptotic structure, and in particular from the assumption of asymptotic predictability (weak cosmic censorship): $\mathscr{I}^{+}\subset\overline{D^{+}(S)}$ where $S$ is a partial Cauchy hypersurface and the closure is in the topology of $\bar{M}$. The reader is referred to hawking73 ; chrusciel01 for a discussion of the reasonability of $\star$. In the regularity result by Chruściel and Costa (chrusciel08, , Thm. 4.11) it is also present though framed again using $\mathscr{I}^{+}$, see also (chrusciel08, , Thm. 4.10). The formulation $\star$ allows typically for shorter and less technical presentations.
As Hawking showed, $\star$ allows one to prove that the expansion is positive on the horizon (hawking73, , Lemma 9.2.2), a fact which ultimately leads to the proof of the second law of black hole thermodynamics which states that the area of black holes is non-decreasing (see chrusciel01 ; minguzzi14d for a proof without smoothness assumptions). The fact that Hawking’s argument on the positivity of $\theta$ can be adapted to the non-smooth case is non-trivial and was proved in (chrusciel01, , Theor. 4.1). These authors still work in the geometry of spacetime admitting a suitable conformal completion. This requires more assumptions, though the framework is compatible with that adopted in this work, in which the horizon is defined via the boundary of $I^{-}(M_{end})$. We adapt the result (chrusciel01, , Theor. 4.1) to the present framework as follows.
Theorem 3.3
@@
A proof can be obtained from the sketch of proof in (minguzzi14d, , Thm. 22) (see also the original reference (chrusciel01, , Thm. 4.1)), with some trivial replacements such as $I^{-}(\mathscr{I}^{+},\bar{M})\to I^{-}(M_{end})$. Of course, the inequality $\theta\geq 0$ is understood in a support function sense, as $H^{+}$ is a priori non-differentiable.
@@
Proof (of Theorem 3.2)
Without loss of generality we can assume that $H^{+}$ is connected, otherwise apply the following argument to a single component (note that the previous results generalize to this case without difficulty).
Under the quotient of Corollary 5 the horizon $H^{+}$ projects to a compact $C^{0}$ future null hypersurface $\hat{H}$ which is topologically a product. Moreover, we have $\theta\geq 0$ on $\hat{H}$, thus by the result in (larsson14, , Thm. 1.43) (minguzzi14d, , Thm. 18) it follows that $\hat{H}$ and hence $H^{+}$ is a totally geodesic lightlike hypersurface as regular as the metric (the Cauchy horizon condition in these theorems is used to infer future completeness of the generators from which $\theta\geq 0$ is inferred, thus for our purposes the Cauchy horizon condition can be replaced by $\theta\geq 0$, indeed (minguzzi14d, , Thm. 13) implies that $\theta=0$, $\mu^{s}_{ij}=0$ and (minguzzi14d, , Thm. 17) implies that it is as regular as the metric). Notice that although through each point of $H^{+}$ passes a unique generator, the generators of $H^{+}$ need not be past-inextendible (in some examples they are past incomplete and can be extended escaping $H^{+}$). $\square$
Remark 2
Summarizing, from asymptotic conditions it is possible to infer condition $\star$, a kind of future completeness condition, from which, in turn, it is possible to obtain, by using the Raychaudhuri equation and the achronality of the horizon, $\theta\geq 0$, and hence the smoothness of the horizon via Theorem 3.2.
One could think of a different strategy for obtaining $\theta\geq 0$. In presence of a Killing field we could try to compactify the horizon, as done in Cor. 5, and apply the dichotomy non-degenerate/degenerate (i.e. all generators are future complete (or all past complete), or all incomplete) of compact horizons moncrief83 ; moncrief08 ; reiris21 ; minguzzi21 . By staying in the future complete case, we could infer $\theta\geq 0$ from the achronality of the horizon and then get the smoothness of the horizon with a proof analogous to the above.
The problem is that available proofs of the mentioned dichotomy apply to compact horizons that are already known to be smooth, so one would have first to extend the ribbon argument and other analytical techniques used in the proof of the dichotomy to the non-smooth horizon setting, a strategy which is not entirely clear could be successfully pursued given the technical difficulties involved.
@@ 4 Existence of cross-sections
Proof
[…]
@@
We recall that a cross-section in the sense of chrusciel08 is a topological submanifold of $H^{+}$ that intersects all generators exactly once. We shall indeed prove that cross-sections exist (in chrusciel08 this was an assumption) by taking advantage of the existence of the trivial principal bundle established in the previous section. Note that contrary to chrusciel08 we do not use the existence of cross-sections to obtain smoothness results, the logical order in this work is reversed ($C^{1}$ regularity is used to obtain cross-sections, see Lemma 6 below).
@@
The next result uses ideas in (friedrich99, , Remark 2.2), see also chrusciel94b . In that remark they used without mention a causal simplicity assumption at the end of their argument (also beware that elsewhere they assume smoothness of the horizon, which, in any case, we proved without using this type of results). We are able to considerably weaken the causality condition.
Lemma 4
@@
Suppose that strongly causality holds on $H^{+}$. Let $r\in M_{end}$ then every Killing orbit of $H^{+}$ intersects $Y:=\partial I^{+}(r)\cap H^{+}$. Moreover, in the forward direction it is bound to enter $I^{+}(r)$ and to remain in it once entered. Similarly, in the backward direction, it is bound to escape $\overline{I^{+}(r)}$ and to remain outside it once escaped. Thus the orbit intersects $Y$ in a compact segment (possibly a point).
@@
If past reflectivity holds in an open set containing $\overline{I^{-}(M_{end})}\cap I^{+}(M_{end})$ then the intersection is just one single point.
@@
The curious and interesting role of past reflectivity in black hole physics has been pointed out in minguzzi20 . We recall that it is implied by causal continuity and hence by global hyperbolicity. It is also implied by the existence of a continuous Lorentzian distance function beem77 (minguzzi18b, , Prop. 5.2). Finally, it is implied by the existence of a timelike Killing field which is complete in the past direction clarke88 (minguzzi18b, , Thm. 4.10). As a consequence, past reflectivity holds on $M_{end}$, however, a priori it might not extend up to the horizon.
@@
It is worth recalling that past reflectivity and past distinction imply stable causality (minguzzi18b, , Thm. 4.111) and that we are imposing strong causality in a neighborhood of $H^{+}$. So imposing the property of past reflectivity implies the existence of a time function in a neighborhood of the horizon.
Proof
@@
Let $p\in H^{+}$ then as $H^{+}\subset I^{+}(\phi(r))$, we can find $\tau>0$ such that $p\in I^{+}(\phi_{-\tau}(r))$ (remember that $k$ is timelike in $M_{end}$ but not necessarily near the horizon), and hence $\phi_{\tau}(p)\in I^{+}(r)$. Moreover, by Prop. 7 we can find $s>0$ such that $p\notin\overline{I^{+}(\phi_{s}(r))}$ thus $\phi_{-s}(p)\notin\overline{I^{+}(r)}$. Thus the Killing orbit $x(t)=\phi_{t}(p)\subset H^{+}$ intersects $\partial I^{+}(r)$ and in the forward direction it enters $I^{+}(r)$ while in the past direction it escapes $\overline{I^{+}(r)}$.
@@
Assume past reflectivity. If $q\in\overline{I^{+}(r)}\cap H^{+}$ then by past reflectivity $r\in\overline{I^{-}(q)}$ and for $\tau>0$, as $r\gg\phi_{-\tau}(r)$, $q\in I^{+}(\phi_{-\tau}(r))$, and by the isometry $\phi_{\tau}(q)\in I^{+}(r)$, namely any orbit can have at most one point in $\partial I^{+}(r)$ as any subsequent point belongs to $I^{+}(r)$. $\square$
Lemma 5
Suppose that strongly causality holds on $H^{+}$ and $S=\pi(H^{+})$ is compact. Let $r\in M_{end}$, then the set $Y:=\partial I^{+}(r)\cap H^{+}$ is non-empty and compact. If past reflectivity holds in a neighborhood of $\overline{I^{-}(M_{end})}\cap I^{+}(M_{end})$, then $Y$ and $S$ are homeomorphic.
Proof